HETAV PATEL

Grade 11,

how to derive unit and dimention of magnetic permeability ?

how to derive unit and dimention of magnetic permeability ?

Grade:12th pass

1 Answers

jitender
6 years ago
Step 1 : write any valid formula which is easiest to you Ex. F = BILF = magnetic force.B= magnetic fieldI= currentL= displacement of initial and final pointsNow B= F/IL _____1st eqnAnd B = ₩Idlsin● / 4^r2 _____2nd eqn₩= magnetic permeabilityFrom 1st and 2nd₩= F4^r2/Idlsin●(IL)Step 2::: remove any numerical unitless constants like 4, ^ sin ₩ = Fr2/Idl (IL)3. To find unit write units ₩ = Nm2/A2m2₩ = N/A2Dimention= MLT-2A-2